OpenStudy (anonymous):

Factor the Expression x^2-15x+54

OpenStudy (unklerhaukus):

can you think of two factors (positive or negative) of 54, that add to -15?

OpenStudy (anonymous):

9 & 6

OpenStudy (unklerhaukus):

right, but we want them to add to negative 15,

OpenStudy (anonymous):

-9 +-6

OpenStudy (unklerhaukus):

yeah, so x^2-15x+54=(x-9)(x-6) now if you want to check it (and you should) expand the brackets in your answer (x-9)(x-6)=x(x-6)-9(x-6) = x^2-6x-9x+54

OpenStudy (unklerhaukus):

which should simplify to the original expression in the question

OpenStudy (anonymous):

thank you! i understand so much better now!

OpenStudy (unklerhaukus):

\[\Large\ddot\smile\]